# tablesection algorithmsection

### v1

@@ 1 Introduction
The concentration of the sample covariance (the sample moment tensor of order $p=2$) has been extensively studied. Most pertinent to our work, the papers [28, 19] identify a notion of effective dimension that determines the sample complexity. In contrast to our results for even order $p>2,$ in the case $p=2$ the sample covariance matrix is known to be minimax optimal under the operator norm (see [22, Theorem 2] and [19, Theorem 4]). Sharp operator-norm concentration inequalities for simple random tensors of arbitrary order were recently established in [2, 12] using the effective dimension from [28, 19]. In this paper, we prove a new sharp concentration inequality in entrywise maximum norm using a notion of effective dimension introduced in [3, 4] to analyze thresholding estimators for structured covariance operators.
Isserlis’s theorem expresses higher-order moments of a Gaussian distribution as functionals of its covariance. Consequently, our problem falls within the general framework of estimating functionals $f(\Sigma)$ of a covariance operator $\Sigma$ [16, 17], where plug-in estimators $f(\widehat{\Sigma})$ based on the sample covariance $\widehat{\Sigma}$ are often severely biased and suboptimal [20, 18]. Bias reduction techniques —such as jackknife, bootstrap, and Taylor expansions— are commonly used to improve efficiency. Our new dimension-free, non-asymptotic error bounds for Isserlis’s estimator show that, in the context of estimating Gaussian moment tensors, plug-in estimation via Isserlis’s formula performs well and strictly outperforms the sample moment estimator.
@@ 2 Preliminaries
$(v_{1}\otimes\cdots\otimes v_{p})_{\ell_{1},\dots,\ell_{p}}:=v_{1,\ell_{1}}\cdots v_{p,\ell_{p}}.$
The corresponding operator norm is given by
@@
$\|T\|_{\max}:=\sup_{v_{k}\in\mathcal{E}_{k},1\leq k\leq p}|\mathopen{}\mathclose{{}\left\langle T,v_{1}\otimes\cdots\otimes v_{p}}\right\rangle|,$
where $\mathcal{E}_{k}:=\{e_{i}\}_{i=1}^{d_{k}}$ is the standard basis of $\mathbb{R}^{d_{k}}.$
@@ 3.1 Symmetric case
$T:=\mathbb{E}X^{\otimes p}:=\mathbb{E}X\otimes\cdots\otimes X,$
@@
$\widehat{T}_{S}:=\frac{1}{N}\sum_{i=1}^{N}X_{i}^{\otimes p}.$
@@
$\big{(}\widehat{T}_{S}\big{)}_{\ell_{1},\ldots,\ell_{p}}:=\frac{1}{N}\sum_{i=1}^{N}X_{i,\ell_{1}}\cdots X_{i,\ell_{p}},$
where $X_{i,\ell}$ denotes the $\ell$-th coordinate of $X_{i}.$ Alternatively, Isserlis’s theorem [15] (also known as Wick’s probability theorem [31]) provides an exact expression for Gaussian moment tensors in terms of second-order moments. Specifically, for any multi-index $(\ell_{1},\dots,\ell_{p}),$
@@
$\big{(}\widehat{T}_{I}\big{)}_{\ell_{1},\ldots,\ell_{p}}:=\sum_{\pi\in\Pi^{2}_{p}}\prod_{(j,k)\in\pi}\widehat{\Sigma}_{\ell_{j},\ell_{k}}.$
@@
Using the notation above, the following bounds hold:
@@
[24, Theorem 3] generalizes Isserlis’s theorem to isotropic distributions. While beyond the scope of this work, we anticipate the advantage of Isserlis’s estimators to extend to that setting. Numerical results in [6] support this claim. ∎
@@ 3.2 Asymmetric case
Let $p$ be an even integer, and let $X=(X^{(1)},\ldots,X^{(p)})\in\mathbb{R}^{d}$ be zero-mean Gaussian random vectors with covariance matrix $\Sigma$, where each block $X^{(k)}\in\mathbb{R}^{d_{k}}$ and $\sum_{k=1}^{p}d_{k}=:d$. For each $k\in\{1,\ldots,p\}$, denote the marginal covariance by $\Sigma^{(k)}:=\mathbb{E}X^{(k)}\otimes X^{(k)}$, and the cross-covariance by $\Sigma^{(j,k)}:=\mathbb{E}X^{(j)}\otimes X^{(k)}$. Let $X_{1},\ldots,X_{N}$ be i.i.d. copies of $X$, where $X_{i}=(X^{(1)}_{i},\ldots,X^{(p)}_{i})$. Our goal is to estimate the moment tensor
$T:=\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)},$
@@
$\widehat{T}_{S}:=\frac{1}{N}\sum_{i=1}^{N}X_{i}^{(1)}\otimes\cdots\otimes X_{i}^{(p)}.$
@@
$\big{(}\widehat{T}_{S}\big{)}_{\ell_{1},\ldots,\ell_{p}}:=\frac{1}{N}\sum_{i=1}^{N}X^{(1)}_{i,\ell_{1}}\cdots X^{(p)}_{i,\ell_{p}},$
where $X^{(k)}_{i,\ell}$ denotes the $\ell$-th coordinate of $X^{(k)}_{i}.$
@@
$\big{(}\widehat{T}_{I}\big{)}_{\ell_{1},\ldots,\ell_{p}}:=\sum_{\pi\in\Pi^{2}_{p}}\prod_{(j,k)\in\pi}\widehat{\Sigma}^{(j,k)}_{\ell_{j},\ell_{k}}.$
@@
Using the notation above, the following bounds hold:
@@
For arbitrary vectors $v_{1},\ldots,v_{p}\in\mathbb{R}^{d}$, we have
$\langle T_{X}-T_{Y},v_{1}\otimes\cdots\otimes v_{p}\rangle=\mathopen{}\mathclose{{}\left\langle\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)}-\mathbb{E}Y^{(1)}\otimes\cdots\otimes Y^{(p)},v_{1}\otimes\cdots\otimes v_{p}}\right\rangle$
@@
[…]
@@
$=\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\bigg{(}\prod_{s=1}^{\ell-1}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{X}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg{)}$
×$\times\!\langle(\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y})v_{\pi(2\ell-1)},v_{\pi(2\ell)}\rangle\bigg{(}\prod_{s=\ell+1}^{p/2}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg{)}.$
To bound the operator norm, we take the supremum over $v_{1},\ldots,v_{p}\in\mathcal{S}^{d}$,
$\|T_{X}-T_{Y}\|=\sup_{v_{1},\ldots,v_{p}\in\mathcal{S}^{d}}|\langle T_{X}-T_{Y},v_{1}\otimes\cdots\otimes v_{p}\rangle|$
$=\sup_{v_{1},\ldots,v_{p}\in\mathcal{S}^{d}}\bigg{|}\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\bigg{(}\prod_{s=1}^{\ell-1}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{X}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg{)}$
×$\times\langle(\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y})v_{\pi(2\ell-1)},v_{\pi(2\ell)}\rangle\times\bigg{(}\prod_{s=\ell+1}^{p/2}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}v_{\pi(2s-1)},v_{\pi(2s)}\bigg{)}\bigg{|}$
$\leq\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\Bigg{[}\sup_{v_{1},\ldots,v_{p}\in\mathcal{S}^{d}}\bigg{|}\bigg{(}\prod_{s=1}^{\ell-1}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{X}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg{)}$
×$\times\langle(\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y})v_{\pi(2\ell-1)},v_{\pi(2\ell)}\rangle\times\bigg{(}\prod_{s=\ell+1}^{p/2}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}v_{\pi(2s-1)},v_{\pi(2s)}\bigg{)}\bigg{|}\Bigg{]}$
$=\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\bigg{(}\prod_{s=1}^{\ell-1}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{X}\|\bigg{)}\|\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y}\|\bigg{(}\prod_{s=\ell+1}^{p/2}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}\|\bigg{)}.$
For $j,k\in\{1,\ldots,p\}$, we introduce the normalized deviation $\varepsilon^{(j,k)}:=\frac{\|\Sigma^{(j,k)}_{X}-\Sigma^{(j,k)}_{Y}\|}{\big{(}\|\Sigma^{(j,j)}_{Y}\|\|\Sigma^{(k,k)}_{Y}\|\big{)}^{1/2}}$ and define $\varepsilon_{*}:=\max_{j\neq k}\varepsilon^{(j,k)}$. Using the inequality $\|\Sigma^{(j,k)}_{Y}\|\leq\|\Sigma^{(j,j)}_{Y}\|^{1/2}\|\Sigma^{(k,k)}_{Y}\|^{1/2}$, we obtain
$\|T_{X}-T_{Y}\|$
$\leq\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\bigg{(}\prod_{s=1}^{\ell-1}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{X}\|\bigg{)}\|\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y}\|\bigg{(}\prod_{s=\ell+1}^{p/2}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}\|\bigg{)}$

### v2

# On the Estimation of Gaussian Moment Tensors

@@ 1 Introduction
The concentration of the sample covariance (the sample moment tensor of order $p=2$) has been extensively studied. Most pertinent to our work, the papers [28, 19] identify a notion of effective dimension that determines the sample complexity. In contrast to our results for even order $p>2,$ in the case $p=2$ the sample covariance matrix is known to be minimax optimal under the operator norm (see [22, Theorem 2] and [19, Theorem 4]). Sharp operator norm concentration inequalities for simple random tensors of arbitrary order were recently established in [2, 12] using the effective dimension from [28, 19]. In this paper, we prove a new sharp concentration inequality in entrywise maximum norm using a notion of effective dimension introduced in [3, 4] to analyze thresholding estimators for structured covariance operators.
[…]
@@ 2 Preliminaries
$(v_{1}\otimes\cdots\otimes v_{p})_{\ell_{1},\dots,\ell_{p}}:=v_{1,\ell_{1}}\cdots v_{p,\ell_{p}},$
where $v_{i,\ell}$ denotes the $\ell$-th coordinate of $v_{i}.$ The operator norm of $T\in\mathbb{R}^{d_{1}\times\cdots\times d_{p}}$ is
@@
$\|T\|_{\max}:=\sup_{v_{k}\in\mathcal{E}^{d_{k}},1\leq k\leq p}|\mathopen{}\mathclose{{\left\langle T,v_{1}\otimes\cdots\otimes v_{p}}}\right\rangle|,$
where $\mathcal{E}^{d_{k}}:=\{e_{i}\}_{i=1}^{d_{k}}$ is the standard basis of $\mathbb{R}^{d_{k}}.$
@@ 3.1 Symmetric case
$T:=\mathbb{E}X^{\otimes p}:=\mathbb{E}X\otimes\cdots\otimes X,$ (3.1)
@@
$\widehat{T}_{S}:=\frac{1}{N}\sum_{i=1}^{N}X_{i}^{\otimes p}.$ (3.2)
@@
$\big(\widehat{T}_{S}\big)_{\ell_{1},\ldots,\ell_{p}}:=\frac{1}{N}\sum_{i=1}^{N}X_{i,\ell_{1}}\cdots X_{i,\ell_{p}}.$
Alternatively, Isserlis’s theorem [15] (also known as Wick’s probability theorem [31]) provides an exact expression for Gaussian moment tensors in terms of second-order moments. Specifically, for any multi-index $(\ell_{1},\dots,\ell_{p}),$
@@
$\big(\widehat{T}_{I}\big)_{\ell_{1},\ldots,\ell_{p}}:=\sum_{\pi\in\Pi^{2}_{p}}\prod_{(j,k)\in\pi}\widehat{\Sigma}_{\ell_{j},\ell_{k}}.$ (3.3)
@@
Let $X\sim\mathcal{N}(0,\Sigma)$ be a zero-mean Gaussian random vector in $\mathbb{R}^{d}$ with covariance matrix $\Sigma,$ and let $X_{1},\ldots,X_{N}$ be i.i.d. copies of $X$. For any even integer $p$, let $T,\widehat{T}_{S},$ and $\widehat{T}_{I}$ be as defined in equations (3.1), (3.2), and (3.3), respectively. Then the following bounds hold.
@@
The upper bounds for the sample moment tensor in Theorem 3.1 hold for sub-Gaussian data. In contrast, our analysis of the Isserlis’s estimator is limited to Gaussian data. An interesting question beyond the scope of this work is to leverage the generalization of Isserlis’s theorem for isotropic distributions in [24, Theorem 3] to define and analyze more general Isserlis’s-type estimators. Numerical results in [6] suggest that the advantage of Isserlis’s-type estimators over the sample moment tensor may carry over to isotropic, sub-Gaussian data. ∎
The following theorem establishes a lower bound for Isserlis’s estimator.
Theorem 3.4.
Let $X\sim\mathcal{N}(0,\Sigma)$ be a zero-mean Gaussian random vector in $\mathbb{R}^{d}$ with covariance matrix $\Sigma,$ and let $X_{1},\ldots,X_{N}$ be i.i.d. copies of $X$. For any even integer $p$, let $T$ and $\widehat{T}_{I}$ be as defined in equations (3.1) and (3.3), respectively. Then the following bounds hold.
(i) Operator norm bound:
$\mathbb{E}\lVert\widehat{T}_{I}-T\rVert\gtrsim_{p}\|\Sigma\|^{p/2}\bigg(\frac{1}{\kappa(\Sigma)^{p/2-1}}\sqrt{\frac{r_{2}(\Sigma)}{N}}+\bigg(\frac{r_{2}(\Sigma)}{N}\bigg)^{p/2}\bigg),$
where $\kappa(\Sigma):=\lambda_{\max}(\Sigma)/\lambda_{\min}(\Sigma)$ is the condition number of $\Sigma$.
(ii) Entrywise maximum norm bound:
$\mathbb{E}\lVert\widehat{T}_{I}-T\rVert_{\max}\gtrsim_{p}\|\Sigma\|_{\max}^{p/2}\bigg(\frac{1}{\kappa(D(\Sigma))^{p/2-1}}\sqrt{\frac{r_{\max}(\Sigma)}{N}}+\bigg(\frac{r_{\max}(\Sigma)}{N}\bigg)^{p/2}\bigg),$
where $D(\Sigma)$ is the diagonal matrix with the same diagonal entries as $\Sigma$, and $\kappa(D(\Sigma))$ denotes the condition number of $D(\Sigma)$.
Proof.
(i) Operator norm bound. By Proposition 3.5 below, we have
$\frac{\|\widehat{T}_{I}-T\|}{\|T\|}\geq\max\mathopen{}\mathclose{{\left\{\frac{\|\widehat{\Sigma}-\Sigma\|}{\|\Sigma\|}\bigg(\frac{1}{\kappa(\Sigma)}\bigg)^{p/2-1},\bigg(\frac{\|\widehat{\Sigma}-\Sigma\|}{\|\Sigma\|}\bigg)^{p/2}}}\right\}.$
Taking expectations on both sides and substituting the bound $\mathbb{E}\|\widehat{\Sigma}-\Sigma\|/\|\Sigma\|\asymp\sqrt{\frac{r_{2}(\Sigma)}{N}}+\frac{r_{2}(\Sigma)}{N}$ from [19, Theorem 4] yields the desired result.
(ii) Entrywise maximum norm bound. By Proposition 3.5 below,
$\frac{\|\widehat{T}_{I}-T\|_{\max}}{\|T\|_{\max}}\geq\max\mathopen{}\mathclose{{\left\{\frac{\|D(\widehat{\Sigma}-\Sigma)\|_{\max}}{\|D(\Sigma)\|_{\max}}\bigg(\frac{1}{\kappa(D(\Sigma))}\bigg)^{p/2-1},\bigg(\frac{\|D(\widehat{\Sigma}-\Sigma)\|_{\max}}{\|D(\Sigma)\|_{\max}}\bigg)^{p/2}}}\right\}.$
A straightforward adaptation of the argument in [12, Proposition 3.1] and [2, Proposition 3.1] —replacing the operator norm with the entrywise maximum norm— yields
$\mathbb{E}\|D(\widehat{\Sigma}-\Sigma)\|_{\max}=\mathbb{E}\sup_{v\in\mathcal{E}^{d}}|\langle(\widehat{\Sigma}-\Sigma)v,v\rangle|\gtrsim\|\Sigma\|_{\max}\bigg(\sqrt{\frac{r_{\max}(\Sigma)}{N}}+\frac{r_{\max}(\Sigma)}{N}\bigg).$
Taking expectations in the previous inequality, substituting this bound, and noting that $\|D(\Sigma)\|_{\max}=\|\Sigma\|_{\max}$ completes the proof. ∎
The following proposition, proved in Section 4, was used in the proof of Theorem 3.4.
Proposition 3.5.
$\frac{\|T_{X}-T_{Y}\|}{\|T_{Y}\|}\geq\max\mathopen{}\mathclose{{\left\{\frac{\|\Sigma_{X}-\Sigma_{Y}\|}{\|\Sigma_{Y}\|}\mathopen{}\mathclose{{\left(\frac{1}{\kappa(\Sigma_{Y})}}}\right)^{p/2-1},\mathopen{}\mathclose{{\left(\frac{\|\Sigma_{X}-\Sigma_{Y}\|}{\|\Sigma_{Y}\|}}}\right)^{p/2}}}\right\},$
where $\kappa(\Sigma_{Y}):=\lambda_{\max}(\Sigma_{Y})/\lambda_{\min}(\Sigma_{Y})$ is the condition number of $\Sigma_{Y}$. Similarly, under the entrywise maximum norm,
$\frac{\|T_{X}-T_{Y}\|_{\max}}{\|T_{Y}\|_{\max}}$ $\geq\max\mathopen{}\mathclose{{\left\{\frac{\|D(\Sigma_{X}-\Sigma_{Y})\|_{\max}}{\|D(\Sigma_{Y})\|_{\max}}\mathopen{}\mathclose{{\left(\frac{1}{\kappa(D(\Sigma_{Y}))}}}\right)^{p/2-1},\mathopen{}\mathclose{{\left(\frac{\|D(\Sigma_{X}-\Sigma_{Y})\|_{\max}}{\|D(\Sigma_{Y})\|_{\max}}}}\right)^{p/2}}}\right\},$
where $D(\Sigma_{X}-\Sigma_{Y})$ and $D(\Sigma_{Y})$ denote the diagonal matrices with the same diagonal entries as $\Sigma_{X}-\Sigma_{Y}$ and $\Sigma_{Y}$, respectively.
@@ 3.2 Asymmetric case
Let $p$ be an even integer, and let $X=(X^{(1)},\ldots,X^{(p)})\in\mathbb{R}^{d}$ be a zero-mean Gaussian random vector with covariance matrix $\Sigma$, where each block $X^{(k)}\in\mathbb{R}^{d_{k}}$ and $\sum_{k=1}^{p}d_{k}=d$. For each $k\in\{1,\ldots,p\}$, denote the marginal covariance by $\Sigma^{(k)}:=\mathbb{E}X^{(k)}\otimes X^{(k)}$, and the cross-covariance by $\Sigma^{(j,k)}:=\mathbb{E}X^{(j)}\otimes X^{(k)}$. Let $X_{1},\ldots,X_{N}$ be i.i.d. copies of $X$, where $X_{i}=(X^{(1)}_{i},\ldots,X^{(p)}_{i})$. Our goal is to estimate the moment tensor
$T:=\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)},$ (3.4)
@@
$\widehat{T}_{S}:=\frac{1}{N}\sum_{i=1}^{N}X_{i}^{(1)}\otimes\cdots\otimes X_{i}^{(p)}.$ (3.5)
@@
$\big(\widehat{T}_{S}\big)_{\ell_{1},\ldots,\ell_{p}}:=\frac{1}{N}\sum_{i=1}^{N}X^{(1)}_{i,\ell_{1}}\cdots X^{(p)}_{i,\ell_{p}}$.
@@
$\big(\widehat{T}_{I}\big)_{\ell_{1},\ldots,\ell_{p}}:=\sum_{\pi\in\Pi^{2}_{p}}\prod_{(j,k)\in\pi}\widehat{\Sigma}^{(j,k)}_{\ell_{j},\ell_{k}}$. (3.6)
@@
Let $p$ be an even integer, and let $X=(X^{(1)},\ldots,X^{(p)})\in\mathbb{R}^{d}$ be a zero-mean Gaussian random vector with covariance matrix $\Sigma$. For each $k\in\{1,\ldots,p\}$, let $\Sigma^{(k)}:=\mathbb{E}X^{(k)}\otimes X^{(k)}$, and for $j,k\in\{1,\ldots,p\}$, let $\Sigma^{(j,k)}:=\mathbb{E}X^{(j)}\otimes X^{(k)}$. Let $X_{1},\ldots,X_{N}$ be i.i.d. copies of $X$, where $X_{i}=(X^{(1)}_{i},\ldots,X^{(p)}_{i})$. Let $T,\widehat{T}_{S},$ and $\widehat{T}_{I}$ be as defined in equations (3.4), (3.5), and (3.6), respectively. Then the following bounds hold.
@@
We begin by observing that
$\|T_{X}-T_{Y}\|$ $=\sup_{v\in\mathcal{S}^{d}}|\mathopen{}\mathclose{{\left\langle\mathbb{E}X^{\otimes p}-\mathbb{E}Y^{\otimes p},v^{\otimes p}}}\right\rangle|$
$=\sup_{v\in\mathcal{S}^{d}}\mathopen{}\mathclose{{\left|\mathbb{E}\langle X,v\rangle^{p}-\mathbb{E}\langle Y,v\rangle^{p}}}\right|$
$=(p-1)!!\sup_{v\in\mathcal{S}^{d}}\mathopen{}\mathclose{{\left|\langle\Sigma_{X}v,v\rangle^{p/2}-\langle\Sigma_{Y}v,v\rangle^{p/2}}}\right|$
$=(p-1)!!\sup_{v\in\mathcal{S}^{d}}\mathopen{}\mathclose{{\left|\langle(\Sigma_{X}-\Sigma_{Y})v,v\rangle}}\right|\,\bigg|\sum_{\ell=0}^{p/2-1}\langle\Sigma_{X}v,v\rangle^{\ell}\langle\Sigma_{Y}v,v\rangle^{p/2-1-\ell}\bigg|,$
where the third equality uses the moment formula for centered Gaussian variables. Taking the supremum over the first term and the infimum over the second yields
$\|T_{X}-T_{Y}\|$ $\geq(p-1)!!\sup_{v\in\mathcal{S}^{d}}\mathopen{}\mathclose{{\left|\langle(\Sigma_{X}-\Sigma_{Y})v,v\rangle}}\right|\cdot\inf_{v\in\mathcal{S}^{d}}\bigg|\sum_{\ell=0}^{p/2-1}\langle\Sigma_{X}v,v\rangle^{\ell}\langle\Sigma_{Y}v,v\rangle^{p/2-1-\ell}\bigg|$
$\geq(p-1)!!\|\Sigma_{X}-\Sigma_{Y}\|\cdot\lambda_{\min}(\Sigma_{Y})^{p/2-1}.$
Since $\|T_{Y}\|=(p-1)!!\,\|\Sigma_{Y}\|^{p/2}$, it follows that
$\frac{\|T_{X}-T_{Y}\|}{\|T_{Y}\|}\geq\frac{\|\Sigma_{X}-\Sigma_{Y}\|}{\|\Sigma_{Y}\|}\cdot\mathopen{}\mathclose{{\left(\frac{1}{\kappa(\Sigma_{Y})}}}\right)^{p/2-1},$ (4.1)
where $\kappa(\Sigma_{Y}):=\lambda_{\max}(\Sigma_{Y})/\lambda_{\min}(\Sigma_{Y})$ is the condition number of $\Sigma_{Y}$.
Next, let $v$ be a unit eigenvector associated with the eigenvalue of $\Sigma_{X}-\Sigma_{Y}$ of largest magnitude, so that $\langle(\Sigma_{X}-\Sigma_{Y})v,v\rangle=\pm\|\Sigma_{X}-\Sigma_{Y}\|$. If the sign is positive, we have
$\|T_{X}-T_{Y}\|$ $\geq(p-1)!!\mathopen{}\mathclose{{\left|\langle\Sigma_{X}v,v\rangle^{p/2}-\langle\Sigma_{Y}v,v\rangle^{p/2}}}\right|$
$=(p-1)!!\mathopen{}\mathclose{{\left|\big(\langle(\Sigma_{X}-\Sigma_{Y})v,v\rangle+\langle\Sigma_{Y}v,v\rangle\big)^{p/2}-\langle\Sigma_{Y}v,v\rangle^{p/2}}}\right|$
$=(p-1)!!\Big(\big(\|\Sigma_{X}-\Sigma_{Y}\|+\langle\Sigma_{Y}v,v\rangle\big)^{p/2}-\langle\Sigma_{Y}v,v\rangle^{p/2}\Big)$
$\geq(p-1)!!\|\Sigma_{X}-\Sigma_{Y}\|^{p/2}.$
If the sign is negative, a symmetric argument gives the same bound with $\Sigma_{X}$ and $\Sigma_{Y}$ interchanged. Therefore,
$\frac{\|T_{X}-T_{Y}\|}{\|T_{Y}\|}\geq\frac{(p-1)!!\|\Sigma_{X}-\Sigma_{Y}\|^{p/2}}{(p-1)!!\|\Sigma_{Y}\|^{p/2}}=\mathopen{}\mathclose{{\left(\frac{\|\Sigma_{X}-\Sigma_{Y}\|}{\|\Sigma_{Y}\|}}}\right)^{p/2}.$ (4.2)
Combining (4.1) and (4.2) yields the desired lower bound in operator norm.
For the entrywise maximum norm, we first obtain a lower bound of the same form as in the operator norm case, except that the supremum is now taken over the standard basis vectors. Specifically,
$\|T_{X}-T_{Y}\|_{\max}$ $=\sup_{v_{k}\in\mathcal{E}^{d},1\leq k\leq p}|\mathopen{}\mathclose{{\left\langle T_{X}-T_{Y},v_{1}\otimes\cdots\otimes v_{p}}}\right\rangle|$
$\geq\sup_{v\in\mathcal{E}^{d}}|\mathopen{}\mathclose{{\left\langle T_{X}-T_{Y},v^{\otimes p}}}\right\rangle|$
$=(p-1)!!\sup_{v\in\mathcal{E}^{d}}\mathopen{}\mathclose{{\left|\langle\Sigma_{X}v,v\rangle^{p/2}-\langle\Sigma_{Y}v,v\rangle^{p/2}}}\right|,$
where $\mathcal{E}^{d}=\{e_{i}\}_{i=1}^{d}$ denotes the standard basis of $\mathbb{R}^{d}$. Then, the desired lower bound for the entrywise maximum norm follows from the same argument as for the operator norm case, with only minor modifications. We omit the details for brevity. ∎
5 Proof of Proposition 3.8
Proof of Proposition 3.8.
For arbitrary vectors $v_{1}\in\mathbb{R}^{d_{1}},\ldots,v_{p}\in\mathbb{R}^{d_{p}}$, we have
$\langle T_{X}-T_{Y},v_{1}\otimes\cdots\otimes v_{p}\rangle$ $=\Big\langle\mathbb{E}X^{(1)}\otimes\cdots\otimes X^{(p)}-\mathbb{E}Y^{(1)}\otimes\cdots\otimes Y^{(p)},v_{1}\otimes\cdots\otimes v_{p}\Big\rangle$
@@
$\overset{\text{(i)}}{=}\sum_{\pi\in\Pi_{p}^{2}}\prod_{(j,k)\in\pi}\mathbb{E}\langle X^{(j)},v_{j}\rangle\langle X^{(k)},v_{k}\rangle-\sum_{\pi\in\Pi_{p}^{2}}\prod_{(j,k)\in\pi}\mathbb{E}\langle Y^{(j)},v_{j}\rangle\langle Y^{(k)},v_{k}\rangle$
@@
$=\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\Bigg[\bigg(\prod_{s=1}^{\ell-1}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{X}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg)$
$\quad\times\langle(\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y})v_{\pi(2\ell-1)},v_{\pi(2\ell)}\rangle\bigg(\prod_{s=\ell+1}^{p/2}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg)\Bigg].$
To bound the operator norm, we take the supremum over $v_{1}\in\mathcal{S}^{d_{1}},\ldots,v_{p}\in\mathcal{S}^{d_{p}}$,
$\|T_{X}-T_{Y}\|=\sup_{v_{k}\in\mathcal{S}^{d_{k}},1\leq k\leq p}|\langle T_{X}-T_{Y},v_{1}\otimes\cdots\otimes v_{p}\rangle|$
$=\sup_{v_{k}\in\mathcal{S}^{d_{k}},1\leq k\leq p}\bigg|\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\bigg(\prod_{s=1}^{\ell-1}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{X}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg)$
$\quad\times\langle(\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y})v_{\pi(2\ell-1)},v_{\pi(2\ell)}\rangle\bigg(\prod_{s=\ell+1}^{p/2}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg)\bigg|$
$\leq\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\Bigg[\sup_{v_{k}\in\mathcal{S}^{d_{k}},1\leq k\leq p}\bigg|\bigg(\prod_{s=1}^{\ell-1}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{X}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg)$
$\quad\times\langle(\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y})v_{\pi(2\ell-1)},v_{\pi(2\ell)}\rangle\bigg(\prod_{s=\ell+1}^{p/2}\langle\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}v_{\pi(2s-1)},v_{\pi(2s)}\rangle\bigg)\bigg|\Bigg]$
$=\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\Bigg[\bigg(\prod_{s=1}^{\ell-1}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{X}\|\bigg)\|\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y}\|\bigg(\prod_{s=\ell+1}^{p/2}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}\|\bigg)\Bigg$].
For $j,k\in\{1,\ldots,p\}$, we introduce the normalized deviation $\varepsilon^{(j,k)}:=\frac{\|\Sigma^{(j,k)}_{X}-\Sigma^{(j,k)}_{Y}\|}{\big(\|\Sigma^{(j,j)}_{Y}\|\|\Sigma^{(k,k)}_{Y}\|\big)^{1/2}}$ and define $\varepsilon_{*}:=\max_{j\neq k}\varepsilon^{(j,k)}$. Using that $\|\Sigma^{(j,k)}_{Y}\|\leq\|\Sigma^{(j,j)}_{Y}\|^{1/2}\|\Sigma^{(k,k)}_{Y}\|^{1/2}$, we obtain
$\|T_{X}-T_{Y}\|\leq\sum_{\pi\in\Pi_{p}^{2}}\sum_{\ell=1}^{p/2}\Bigg[\bigg(\prod_{s=1}^{\ell-1}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{X}\|\bigg)$
$\quad\times\|\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{X}-\Sigma^{(\pi(2\ell-1),\pi(2\ell))}_{Y}\|\bigg(\prod_{s=\ell+1}^{p/2}\|\Sigma^{(\pi(2s-1),\pi(2s))}_{Y}\|\bigg)\Bigg$]